OpenStudy (anonymous):

Help please!!! Find F (0) for the function f (x)=e 3x+2. Round to the nearest ten-thousandth.

OpenStudy (anonymous):

am lost

OpenStudy (anonymous):

thats the link to the problem

OpenStudy (anonymous):

That was an incomplete response. My computer froze. f(0) = f(x) when x = 0. So to evaluate f(0) for any function, f(x), simply plug in x = 0. That holds true for whatever is "taking the place of the x." So if it said f(7), you'd plug in 7. Or for f(1.56), you'd plug in 1.56.

OpenStudy (anonymous):

sorry computer froze had to reboot also

OpenStudy (anonymous):

so my answer would be 2.00

OpenStudy (anonymous):

Could you show what you did? That's not what I got. I'm assuming the (3x + 2) after the e is an exponent?

OpenStudy (anonymous):

am not sure what i did

OpenStudy (anonymous):

should it be 6x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Plug in x = 0 and solve. e^(3(0) + 2) = e^(0 + 2) = e^2...

OpenStudy (anonymous):

wow I dont understand

OpenStudy (anonymous):

Alright, first, just look at the exponent. Pretend for a second that the problem is f(0) = 3x + 2 and it asks you to solve. Can you solve 3x + 2 when x = 0?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

0+2

OpenStudy (anonymous):

So, the answer is 2, yes?

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

yes 2

OpenStudy (anonymous):

So you can figure out that f(0) = 3x + 2 = 2. Now, it's the same thing, but you have one more step. Instead of the problem being 3x + 2, it's f(x) = e^3x+2. So solve (3x+2) when x = 0, then that number is your exponent for e.

OpenStudy (anonymous):

3(2)+6=12

OpenStudy (anonymous):

I don't understand where this 6 is coming from. The equation is e^3x+2. Plug in 0 = x and solve. It would help me more to help you if you showed your work so I can see where you're going wrong.

OpenStudy (anonymous):

3x2

OpenStudy (anonymous):

Why are you multiply 3 and 2? It's not 3 times 2, the "X" is a variable. It doesn't stand for multiplication, but "3 times x." X is an unknown.

OpenStudy (anonymous):

oh that is what i thought i had to do what take the 3 times the 2 because i thought i had to use it to the power and thats where the 6 came in

OpenStudy (anonymous):

No. It's 3x + 2, and the problem gives you x = 0. So plug in 0 = x, and solve. Then take that and raise e to that power.

OpenStudy (anonymous):

so then if i did that I would take 3x+5= 3(0)+2= 5

OpenStudy (anonymous):

3 * 0 isn't 3. Anything multiplied by 0 is 0.

OpenStudy (anonymous):

yes i know that so that would give me 2

OpenStudy (anonymous):

Correct. but that isn't your final step. You still have e^2. So raise e to the second power, then round according to what the problem asks of you, and voila. You have it.

OpenStudy (anonymous):

You solved for the exponent, but don't forget the e. You still have to do that part.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

7.38

OpenStudy (anonymous):

You question requires that you round to the nearest ten-thousandth place. That is the hundredth place.

OpenStudy (anonymous):

7.3800

OpenStudy (anonymous):

What is e^2 to 7 decimal places?

OpenStudy (anonymous):

I believe it would be 7.4

OpenStudy (anonymous):

sorry I know I am frustrating you just have not done this kind of math before

OpenStudy (anonymous):

You're not frustrating me. It's ok. The best way for me to help you is if you go step-by-step. I want you to write out the answer to e^2 to 7 decimal places.

OpenStudy (anonymous):

0000074

OpenStudy (anonymous):

Not quite, no. I assume to get e^2, you are using a calculator? Write out EXACTLY what the calculator says when you input e^2.

OpenStudy (anonymous):

ok when i input e^2 in the calculator i get 7.38905609893065

OpenStudy (anonymous):

Good. Now the ten thousandth place is 4 places after the decimal. To round off to the ten-thousandths place, round that decimal to 4 places.

OpenStudy (anonymous):

that would be where the 0 is correct?

OpenStudy (anonymous):

Yes. That is the 4th place after the decimal. But remember your rounding rules...

OpenStudy (anonymous):

7.3891

OpenStudy (anonymous):

Yes. Correct.

OpenStudy (anonymous):

thank u. I am so sorry for all the trouble

OpenStudy (anonymous):

No trouble at all. As long as you understand what you just did.

OpenStudy (anonymous):

yes finally

OpenStudy (anonymous):

Good